OpenStudy (anonymous):

In a classroom, 5 out 13 students surveyed said they had two or more siblings. If the entire school has 520 students, predict the number of students that have two or more siblings. Write a proportion to model the problem.

OpenStudy (anonymous):

make a proportionate fraction of 5/13 with 520 being the denominator (5*40)/(13*40) = 200/520 ~ 200 students have two or more siblings.

OpenStudy (anonymous):

\[\frac{5}{13} \times 520 = 5 \times 40 = 200\]